In 10 years a child will be 3 years older than twice her current age. What is her current age
Answer by checkley79(3341)

X+10=3+2X
X-2X=-10+3
-X=-7
X=7 ANS. FOR THE CURRENT AGE.
PROOF:
7+10=3+2*7
17=3+14
17=17